OpenStudy (anonymous):

Rationalize √10+√5 ________ √10-√5

OpenStudy (anonymous):

(a-b)(a+b) = a^2 -b^2 Use that fact and the face that squaring a square root undoes the square root. \[\sqrt(a)^2 = a\]

OpenStudy (anonymous):